OpenStudy (chey):

p+9/5=q *show work please

OpenStudy (anonymous):

solve for?

OpenStudy (anonymous):

p=q-9/5

OpenStudy (chey):

solve for q

OpenStudy (anonymous):

p+9/5=q

OpenStudy (chey):

can i get work so i can understand dat.?

OpenStudy (anonymous):

there is no work the question is complete

OpenStudy (chey):

\[p+9\div5=q\]

OpenStudy (anonymous):

make a table of values for each equation. then graph y=/x/+4